### v1

# Critical behavior and multistability in quasi-two-dimensional turbulence

Abstract
Two-dimensional (2D) turbulence, despite being an idealization of real flows, is of fundamental interest as a model of the spontaneous emergence of order from chaotic flows. The emergence of order often displays critical behavior, whose study is hindered by the long spatial and temporal scales involved. Here, we experimentally study turbulence in periodically driven nanofluidic channels with a high aspect ratio using superfluid helium. We find a multistable transition behavior resulting from cascading bifurcations of large-scale vorticity and critical behavior at the transition to quasi-2D turbulence consistent with phase transitions in periodically driven many-body systems. We demonstrate that quasi-2D turbulent systems can undergo an abrupt change in response to a small change in a control parameter, consistent with predictions for large-scale atmospheric or oceanic flows.
Chaotic, turbulent flows are ubiquitous and occur in vastly diverse systems ranging from a flow around a swimming penguin Masud et al. (2022) to climate systems on Earth Boffetta and Ecke (2012). In homogeneous isotropic turbulence in three dimensions, kinetic energy is transported to progressively smaller scales until dissipated by viscosity, thus destroying any large-scale features Davidson (2015). However, confined to two dimensions, turbulence transports energy to large scales in an inverse cascade, which can result in large-scale self-organization Boffetta and Ecke (2012).
The self-organized state is not necessarily unique, and sudden changes between distinct metastable states are possible. Bifurcations in quasi-2D flows are of particular relevance to the dynamics of geophysical flows. For example, large-scale circulating flow in the Atlantic Ocean is likely to transition into a new regime with a significantly altered global climate Ditlevsen and Ditlevsen (2023); van Westen et al. (2024). More detailed simulations suggest a rich landscape of multiple stable states of large geophysical flows Lohmann et al. (2024). Numerical models of idealized 2D turbulence similarly show multiple stable states, such as large vortices and zonal jets between which the flow transitions stochastically Bouchet and Simonnet (2009), although these transitions are rare Bouchet et al. (2019). There is a growing body of evidence, however, that turbulent flows display analogous behavior to phase transitions in condensed matter systems Cortet et al. (2010) or various nonequilibrium phase transitions (e.g., directed percolation) Sano and Tamai (2016); Hof (2023).
Critical behavior in quasi-2D flows is unexpectedly rich and remains poorly understood Alexakis (2023). Therefore, in this work, we study driven oscillatory turbulent flow confined to channels of aspect ratio $2000D\times 2000D\times D$, $D\approx 500$ nm. We experimentally show critical behavior consistent with random organization Corté et al. (2008) and that the turbulence develops through a cascade of saddle-node bifurcations when the large-scale friction is sufficiently low, and the forcing length scale is sufficiently small.
Flows of superfluid 4He are a superposition of viscous normal flow and inviscid superflow Tilley and Tilley (1990). Microscopically, the superfluid flow is irrotational Tilley and Tilley (1990). Nonzero circulation can exist around topological defects in the superfluid, quantized vortices, which carry a single quantum of circulation $\kappa\approx 9.97\times 10^{-8}$ m2s-1. Quantized vortices facilitate friction-like coupling (mutual friction) of superfluid and normal fluid flows Barenghi et al. (1983). On scales larger than the intervortex spacing, the flow of superfluid helium can be described using coupled equations for superflow and normal fluid flow Donnelly (1991), which in our case reduce to $\mathbf{v}_{n}\approx 0$ (due to viscous clamping Rojas and Davis (2015); Souris et al. (2017)) for normal velocity and for superfluid velocity $\mathbf{v}_{s}$ Donnelly (1991)
$\rho_{s}\frac{\mathrm{D}\mathbf{v}_{s}}{\mathrm{D}t}=-\frac{\rho_{s}}{\rho}\nabla p-\rho_{s}\alpha\kappa L\mathbf{v}_{s}+\rho_{s}\alpha^{\prime}\omega_{s}\times\mathbf{v}_{s}+\mathbf{f},$ (1)
[…]
We study two-dimensional flows using nanofluidic Helmholtz resonators, shown in Fig. 1(a), which consist of a circular volume (“basin”) connected to the surrounding helium bath at saturated vapor pressure by two channels. Similar devices have previously been used for the study of 3He superfluidity Shook et al. (2020, 2024), finite-size effects in 4He Varga et al. (2022), quasi-two-dimensional turbulence Varga et al. (2020) and superfluid optomechanics Spence et al. (2023). The superfluid helium in the channels oscillates under the restoring force provided by the compressibility of the helium and the deformation of the quartz substrate Rojas and Davis (2015). The top and bottom walls of the nanofluidic volume are coated with a 50 nm aluminum layer used for capacitive actuation and sensing (see appendix Supplementary Information and Rojas and Davis (2015); Souris et al. (2017) for details). We studied three flow geometries with different energy injection length scales, whose velocity profile of the Helmholtz mode is shown in Fig. 1(a) — a flow around a sharp corner (device C); a constriction producing an oscillating jet (device J) and a flow through a regular grid (device G; see appendix Supplementary Information for exact dimensions and data from additional device similar to C with rounded corners). Height $D\approx 500$ nm was used for all geometries.
[…]
The high aspect ratio of the flow indicates that the turbulence will be two-dimensional Benavides and Alexakis (2017); Ecke (2017). However, in superfluid helium, the core of the quantized vortex sets the smallest length scale, approximately 0.15 nm Tilley and Tilley (1990). Numerical simulations based on the vortex filament model (see appendix Supplementary Information Sec. 6.2) show that the lifetime of intrinsically three-dimensional vortex configurations is significantly shorter than the flow oscillation period. They are thus transient and cannot participate in long-lived turbulent structures.
Figure 1: Nanofluidic devices and their response. (a) Typical dimensions of the nanofluidic Helmholtz resonators. Dark-shaded areas show aluminum electrodes deposited on the top and bottom walls. The 0.5 mm quartz substrates are not shown. The entire device is submerged in superfluid 4He. A finite element simulation shows the velocity profile of the used acoustic mode. (b) Typical resonant response of the device normalized by the drive. (c) Peak velocity as a function of the driving amplitude of device C at 1.45 K, 1.65 K, and 1.85 K (top to bottom); shaded areas indicated the excess nonlinear damping.
To observe the development of turbulence, the Helmholtz resonance was excited on its resonance frequency with slowly swept drive amplitude in increasing and decreasing directions. The typical response showing the excess dissipation past a critical velocity is shown in Fig. 1(c). Since the flow can transition stochastically between several states, the drive sweep is repeated in several cycles to fully sample the possible states. The vortex density calculated from the excess damping (appendix Supplementary Information (S17)) is shown in Fig. 2(a–i) as two-dimensional histograms for three geometries for a high, intermediate, and low temperature (1.85 K, 1.65 K and 1.45 K, respectively), which determines the large-scale friction in (1). Data from all devices and temperatures are shown in appendix Supplementary Information, Figs. S11–S14.
At high temperatures above a critical velocity, the observed vortex density gradually and reversibly increases above zero for all geometries. For lower temperatures, the behavior strongly depends on geometry. For device C (panels (a–c)), weak hysteresis is observed after the first critical velocity. For device J (panels (d–f)), the flow transitions to a regime with increased $L$ through a discontinuity with stochastic switching between two discrete states. For device G (panels (g–i)), we observe several bifurcations below approximately 1.65 K, resulting in multiple metastable states and hysteresis in the transition region. Additionally, at least two intermittently switching transitions are observable in panel (g). In Fig. 2(j), we show the bifurcation diagram of the branching of the vortex density as a function of temperature. Each branch separated by discontinuous transitions from the others is characterized by a dimensionless quantity $\langle L\kappa^{2}/v^{2}\rangle$. For device J, two branches are observed from 1.85 K to 1.4 K, where the upper branch divides into two. For device G, the branches proliferate in a fractal fashion below 1.45 K. However, the growth in complexity results in a less reliable classification of the data into separate branches.
Figure 2: Vortex density bifurcation diagrams. (a–i) Two-dimensional histograms of vortex density and velocity (darker color means higher probability; logarithmic color scale) for device type C (panels a–c), J (d–f) and G (g–i) for three temperatures, 1.45 K (a, d, g), 1.65 K (b, e, h), and 1.85 K (c, f, i). (j) The temperature-density bifurcation diagram shows the average value of $L\kappa^{2}/v^{2}$ calculated for each individual branch of $L(v)$. Red, device J, black, device G. For $T<1.45$ K, the data for device G continue to branch in a fractal manner (see appendix Supplementary Information, Fig. S13).
To describe the appearance of the discontinuous transitions in devices J and G, we seek a low-dimensional model of the vortex density $L$ (similarly to ref. Varga et al. (2020)). The behavior can be qualitatively captured with a dynamical system (for dimensionless quantities)
$\frac{\mathrm{d}L_{j}}{\mathrm{d}t}=-d_{l}L_{j}-\left(\frac{L_{j}}{1+L_{j}^{2}}\right)^{2}+g_{j},$ (2)
where $L_{j}$ is the vortex density in mode $j$ (i.e., a wave vector for a spatially infinite or periodic system), $j=1...N$, where $N$ is the number of participating modes; $d_{l}$ represents the strength of the large-scale damping due to mutual friction (first term) relative to the pair annihilation of colliding vortices (second term); and $g_{j}$ represents the forcing, which, in general, will be a nonlinear function of the flow velocity and geometry. For the total vortex density, we take $L^{2}=\sum_{j}L^{2}_{j}$. The essential feature of the model is that, for large populations $L>1$, the annihilation of vortices is suppressed. A typical curve $\mathrm{d}L/\mathrm{d}t$ for $N=1$ is shown in Fig. 3(a), and a bifurcation diagram with drive $g$ as the control parameter is shown in Fig. 3(b). For low damping $d_{l}$, the time derivative exhibits a local minimum, resulting in three fixed points, two stable and one unstable, for a certain range of $g$, which qualitatively describes the bifurcation seen in device J. With increasing temperature, the mutual friction increases, increasing large-scale friction on the vortex system. Indeed, the bistability is lost when the parameter $d_{l}$ is increased, similarly to the experiment (cf. Fig. 3(b) and Fig. 2 middle row).
The energy injection scale $l_{f}$ for the device J is close to the width of the constriction, about half of the maximum size of the system $l_{0}$. Therefore, only a small number of wave vectors can exist in the inverse cascade between $l_{f}$ and $l_{0}$. In device G, we increase the ratio $l_{0}/l_{f}$, thus extending the number of modes participating in the inverse cascade. Therefore, we set $N>1$ and assume that the forcing is localized at $j=1$, $g_{1}=g$ and that the higher modes are forced in a cascading manner $g_{j}=cL_{j-1}$, $j>1$, where $c$ characterizes the mode coupling ($d_{l}$ and $c$ are kept constant for simplicity). For $d_{l}=0.05$, $c=0.04$, the total density $L$ in Fig. 3(c) shows multistability due to overlapping pairs of saddle-node bifurcations resembling the experimental data of device G (full lines show stable stationary solutions; dashed lines the unstable fixed points). Although the coupling between vorticity modes is simplified Kraichnan and Montgomery (1980), multistable behavior can occur for more complicated multiplicative coupling Huang et al. (2016). Extending the interpretation of the single-mode dynamics of device J, the resulting physical picture is that, as a particular mode is sufficiently excited, the vortices cluster in a certain pattern to reduce pair annihilation. The energy in the mode thus increases, which allows nonlinear interactions to force progressively larger scales until the size of the system is reached.
The suppression of the annihilation is responsible for the appearance of a pair of bifurcations and the bistable regime. The exact form of the annihilation term does not affect the qualitative behavior of the solutions as long as a local minimum appears. The rational annihilation term in (2) was chosen for simplicity and to maintain the expected $L^{2}$ behavior for low $L$ corresponding to the annihilation of randomly distributed vortices Varga and Skrbek (2018). Since the vortices are expected to be approximately point-like (see appendix Supplementary Information Fig. S7), suppression of annihilation at high densities can occur due to spatial clustering of like-signed vortices, suggesting the emergence of order.
Note that the bistability observed here is different than observed previously in 1 µm confinement Varga et al. (2020), where the bistability was long-lived on the scale of the entire duration of the experiment (i.e., minutes), whereas here we observe intermittent switching or pronounced multistability. However, present results continue the trend observed in Varga et al. (2020) of weakened bistability for C-type device as $D$ decreases. Both types of multi-stability should be considered as special cases of more general dynamical behavior, a complete understanding of which is not yet available.
Figure 3: Nonlinear dynamical model of quasi-2D turbulence. (a) The single-mode model 2 for $N=1$, $d_{l}=$ 0.05, 0.1, and 0.15 and $g=0$, $g>0$ shifts the curve along the y-axis. (b) Stable (full lines) and unstable (dashed lines) stationary solutions of (2) as a function of $g$ (normalized by the position of the local minimum or the inflection point $g_{c}$ of (2)) for the same values of $d_{l}$ as in panel (a). (c) Multistability with several overlapping pairs of saddle-node bifurcations develops by coupling multiple modes of vortex density in a cascading manner as in (2), $N=4$. The orange dashed lines show unstable stationary solutions.
To better understand the intermittent region in device J, we recorded a 32-second-long time series of the response amplitude (an example is shown in appendix Supplementary Information, Fig. S17) for 1.45 K, 1.65 K, and 1.85 K at a set of fixed drive amplitudes. Since the timescale of the forcing is set by the flow oscillation period (about 0.5 ms, see Fig. 1(b)), even relatively short datasets allow the study of the long-time behavior of the flow. The resulting two-dimensional histogram in Fig. 4(a) shows a bimodal shape equivalent to the ramped drive data in Fig. 2. The observed time series was threshold-binarized to detect transitions which were then used to calculate the lifetimes of the more dissipative (lower velocity) state as the average time between the system jumping into and out of the state. The probability density function of the state lifetime is exponential (see appendix Supplementary Information Fig. S16), suggesting that the transitions are memoryless. The mean lifetime $\tau$ as a function of the velocity, shown in Fig. 4(b), shows a diverging behavior, which can be captured by a power law of the form $\tau\propto(v_{c}-v)^{-\eta}$ with $\eta\approx 1.33$ (inset of Fig. 4(b)).
Figure 4: Critical behavior of turbulent transition. (a) Histograms of observed velocity amplitude for a fixed drive at 1.45 K, 1.65 K, and 1.85 K (left to right) for device J. The rug plot indicates the used drives. (b) The mean lifetime of the more dissipative (lower velocity) state as a function of velocity in that state for the three temperatures shifted by the critical velocity (indicated in the histograms by horizontal dashed lines). (inset) Lifetime against reduced dimensionless velocity $(v_{c}-v)/v_{c}$ in logarithmic scale. The black dashed line indicates a power-law with a critical exponent of 1.33.
The appearance of the fluctuating regime in device J resembles reversible-irreversible transitions in periodically driven many-body systems Reichhardt et al. (2023). In particular, the time to reach a steady state in periodically driven systems exhibiting random organization (i.e., a dynamical phase transition into a self-organized state through chaotic dynamics) displays the same critical exponent Corté et al. (2008) as the lifetimes of the turbulent states in Fig. 4. The exponent is also close to the longitudinal correlation critical exponent of 2+1D directed percolation $\xi_{||}\approx 1.295$ Avila et al. (2011); Manneville (2016); Hof (2023). The essential feature of our dynamical model (2) describing the formation of multistability is the decrease in the rate of vortex annihilation at high densities, which is consistent with the spatial separation of vortices of opposite signs, that is, the formation of a stable pattern, possibly driven by the inverse turbulent cascade. Similar clustering was observed in the transient evolution of few-vortex systems in Bose-Einstein condensates Gauthier et al. (2019), superfluid helium Sachkou et al. (2019), and numerically in systems with odd viscosity de Wit et al. (2024), although the dynamics of this clustering in steady-state driven turbulence remain numerically largely unexplored.
Our results show that periodically driven quasi-2D flows are closely related to a broad class of systems displaying reversible-irreversible transitions Reichhardt et al. (2023) despite being turbulent and with a fluctuating number of constituent vortex points. The aspect ratios of the studied flows are comparable to geophysical flows, and thus, the presented results suggest the possibility that similar self-organizing behavior occurs in large-scale planetary flows, such as the hexagonal structure in the polar region of Saturn Yadav and Bloxham (2020), or in the atmosphere of Jupiter Marcus (2004) and will follow similar critical behavior as shown in Fig. 4. Indeed, self-organization due to the inverse cascade was recently shown numerically for a dry atmosphere Alexakis et al. (2024). The observed saddle-node bifurcations give experimental support to the recently raised possibility of the collapse of the Atlantic meridional overturning circulation Ditlevsen and Ditlevsen (2023); van Westen et al. (2024) with significant consequences to the global climate.
Acknowledgements.
We would like to thank K. Olejník for providing access to the clean rooms at FZU and K. Uhlířová for AFM characterization. The authors also thank L. Skrbek, P. Švančara and R. Dwivedi for fruitful discussions. The work was supported by Charles University under PRIMUS/23/SCI/017. CzechNanoLab projects LM2023051 and LNSM-LNSpin funded by MEYS CR are gratefully acknowledged for the financial support of the sample fabrication at CEITEC Nano Research Infrastructure and LNSM at FZU AVČR. AFM characterization was performed in MGML laboratories supported by the program of Czech Research Infrastructures (project no. LM2023065).
Supplementary Information
This supplemental information contains additional details on the characterization of nanofluidic devices, derivation of the relationship between vortex density and fourth sound damping, numerical simulations related to confined superflow, and plots showing the complete dataset whose subset was shown in Fig. 2 in the main text.
Appendix A Nanofluidic Helmhholtz resonators
The flows studied are realized in nanofluidic Helmholtz resonators similar to those used in refs. Rojas and Davis (2015); Souris et al. (2017); Shook et al. (2020); Varga and Davis (2021); Varga et al. (2022). A semi-open nanofluidic channel is etched onto a fused silica substrate, two of which are then bonded together to form a fully enclosed nanofluidic cavity approximately 500 nm in height (see Rojas and Davis (2015); Souris et al. (2017) for fabrication details). The cavity consists of a circular basin (5 mm in diameter) and two channels connecting the basin with the bulk 4He (see the schematic of a resonator in Fig. S1(a)). The characteristic length scale at which a viscous fluid follows the motion of an oscillating wall is the viscous penetration depth $\delta=\sqrt{2\nu/\omega}$, where $\nu$ is the kinematic viscosity and $\omega$ the flow oscillation frequency. For the typical resonance frequency 2 kHz and temperature 1.65 K, and using the normal fluid kinematic viscosity from Donnelly and Barenghi (1998), $\delta\approx$2.7 $\mu$m, which is 5-times greater than the spacing between the upper and lower walls of the cavity. We can, therefore, consider the viscous normal fluid component of superfluid helium to be at rest. Four different chip geometries were used in the study, shown in Fig. S2, which provide different velocity profiles of the superfluid flow inside the channel. The geometry was designed to obtain similar Helmholtz resonance frequencies.
Fig.  S1: Nanofluidic resonator and its response. (a) Sketch of the Helmholtz resonator with typical dimensions. The transparent gray represents the cavity filled with superfluid 4He. The aluminum electrodes cover the entire top and bottom wall of the nanofluidic cavity with a 10 µm gap between the side walls and the edge of the electrode. The substrate enclosing the nanofluidic volume is not shown. (b) A photograph of an assembled Helmholtz resonator used for measurement next to a 5 cent Euro coin for scale. (c) The capacitance bridge circuit used for the drive and detection of the flow. The center-tapped transformer excites both sides of the bridge circuit with an alternating voltage with $\pi$ phase shift between the bridge arms. When the bridge is tuned, only the current resulting from the oscillation of the device capacitance flows through the lock-in amplifier. In order to avoid measuring at the second harmonic (since the electrostatic force is proportional to voltage squared), the device is biased by (typically) a 10 V DC voltage. (d) Total detected current magnitude against frequency for several driving voltage amplitudes. For a sufficiently strong drive, the peak becomes non-Lorentzian with a flattened top.
The flow in the resonator is driven by two aluminum electrodes deposited opposite each other inside the cavity. An applied alternating voltage causes deformation of the substrate, leading to pressure oscillation inside the cavity and, thus, to a pressure gradient along the channels. The resulting motion of the superfluid causes a fluctuation in the capacitance of the device $\delta C(t)$. This fluctuating capacitance is detected using a bridge circuit shown in Fig. S1(c). The capacitance of the lower branch $C$ is adjusted to the static capacitance of the device $C_{0}$ (typically about 400 pF). Sufficiently far from the Helmholtz resonance, zero current is measured by the lock-in amplifier when the bridge is tuned. As the capacitance fluctuations increase close to the resonance, the charge on the device starts to oscillate due to the bias, which is detected by the lock-in amplifier as an oscillating current.
Fig.  S2: Geometry of the resonator channels. The height of all four cavities is 500 nm. The aluminum electrodes are not shown.
The frequency $f_{0}$ of the Helmholtz resonance is given by Souris et al. (2017)
$f_{0}=\dfrac{1}{2\pi}\sqrt{c\dfrac{\rho_{s}ak_{p}}{l\rho^{2}A^{2}(1+\Sigma)}},$ (S1)
where $\Sigma=(\chi Dk_{p})/(2A)$, $A$ is the area of the basin, $a$ the cross-sectional area of the channel and $l$ its length, and $\chi$, $\rho$ and $\rho_{s}$ are isothermal compressibility, and total and superfluid densities. Compresibility values $\chi$ are taken from Brooks and Donnelly (1977), the total density $\rho$ and the superfluid density $\rho_{s}$ from Donnelly and Barenghi (1998). The parameter $c$ is one of the geometric correction parameters introduced by Souris et al. Souris et al. (2017) that corrects for the finite flow in the basin. The measured resonance frequencies for the four devices, along with a fit to (S1), are shown in Fig. S3. Furthermore, we find an improvement of the fit if we allow for a small systematic error in the temperature calibration as $T_{r}=bT$, where $T$ is the measured temperature, $T_{r}$ is the actual temperature, and $b$ is the correction parameter. The correction parameters are shown in Tab. 1.
Fig.  S3: (a) Resonance frequencies for temperatures between 1.325 K and 1.950 K for the four used chip geometries. The dashed line is fit to (S1) considering the temperature correction; for the fit parameters, see Tab. 1. (b) The corresponding widths (FWHM) of the resonances. The width is equal to the linear part of the damping of the resonance motion. The black dashed line is a fit of the sum of (S26) and (S27).
Chip Type $b$ $c$ $\lambda$
Corner 0.97 1.77 1.94 $\pm$ 0.08
Rounded corner 0.97 2.54 2.05 $\pm$ 0.07
Jet 0.97 1.82 2.07 $\pm$ 0.10
Grid 0.97 2.78 1.77 $\pm$ 0.09
Table 1: Values of correction parameters obtained by fitting Eq. (S1) in case of $b$ and $c$, and the sum of Eq. (S26) and (S27) in the case of $\lambda$. The error of $b$ correction factor is $\pm$ 10-3 and of $c$ correction factor $\pm$ 2$\times$10-3.
Appendix B Surface roughness and the first critical velocity
Due to the stringent requirements of the surface cleanliness for the bonding of the resonators, any manipulation of the unbonded chips outside of the cleanroom environment is essentially destructive, therefore we were unable to measure the topography on the actual devices used in the study. However, a scan of the aluminum layer surface topography using atomic force microscopy is shown in Fig. S4 on devices created using a nearly identical fabrication process. The scan shows 10 µm square on the edge of the aluminum. The step height is 48.64 $\pm$ 0.04 nm, and the RMS surface roughness (that is, the root-mean-square distance to the mean plane) of the aluminum layer is approximately 4.1 nm.
Due to the small size of the vortex core, $a_{0}\approx 0.15$ nm, vortices will be pinned on surface roughness and will be unable to slide unless subject to superflow faster than the depinning velocity, which was estimated by Schwarz Schwarz (1985) for a semi-spherical protuberance of radius $b$ on a planar surface as
$v_{\mathrm{pin}}=\frac{\kappa}{2\pi D}\ln\left(\frac{b}{a_{0}}\right).$ (S2)
Replacing the protuberance size $b$ with the measured surface roughness results for $D=500$ nm in $v_{\mathrm{pin}}\approx$ 11 cm/s, which is within about factor 2 of the initial reversible transitions to nonlinear dissipation seen in Fig. S10 and Figs. S11S14. The origin of the slight temperature dependence of this critical velocity is unknown. However, (S2) was obtained for steady flow; in the present case, the temperature dependence is probably the result of the interaction of the temperature-dependent flow oscillation frequency and the vortex relaxation time given by the mutual friction.
Fig.  S4: Atomic force microscopy scan of the topography of the edge of the aluminum surface (higher, rougher area) in the basin area.
Appendix C Helmholtz resonance and vortex density
Similarly to the second sound Varga et al. (2019), the fourth sound used in this work is damped by the presence of quantized vortices, and its damping can be used to calculate the average number of vortices per unit area (vortex density $L$). For completeness, we briefly reproduce the derivation of the equation of motion of the Helmholtz mode and the damping due to quantized vortices based on refs. Varga et al. (2020); Novotný et al. (2024).
Since only the superfluid component is moving inside the channels, the change in the density of the fluid within the basin can be expressed as
$\delta\rho=\delta\left(\dfrac{M_{B}}{V_{B}}\right)=\dfrac{\delta M_{B}}{V_{B}}-\dfrac{M_{B}}{V_{B}^{2}}\delta V_{B}=\dfrac{1}{V_{B}}(-2a\rho_{s}y+2\rho Ax),$ (S3)
where $y$ is the displacement of the superfluid in the channels (positive for outflow), $x$ is the average displacement of the basin walls (positive for compression), the fluid mass $M_{B}=\rho V_{B}$, the volume of the basin $V_{B}=AD$ ($A$ is the area of the basin and $D$ its height). Changes in density and pressure are connected by isothermal compressibility $\delta\rho=\rho\chi\delta P$ and thus we can write
$\delta P=\dfrac{1}{\rho\chi V_{B}}(2\rho Ax-2a\rho_{s}y).$ (S4)
The electrostatic force acting between the charged basin plates $F=C_{0}U^{2}/(2D)$ ($C_{0}$ is the static capacitance of the chip and $U$ the total applied voltage) is equal to contributions from plate deformation and pressure
$F=k_{p}x+A\delta P,$ (S5)
where $k_{p}$ is the spring constant and we neglect the inertia of the basin walls. Combining (S3),(S4),(S5), we write the equation of motion for the fluid displacement $y$ in the channel with length $l$
$\rho_{s}al\ddot{y}$ $=\dfrac{\rho_{s}}{\rho}a\delta P-F_{f},$ (S6)
$\rho_{s}al\ddot{y}$ $=\dfrac{\rho_{s}ak_{p}}{\rho^{2}(\chi V_{B}k_{p}+2A^{2})}\left[\dfrac{2A\rho}{k_{p}}F-2a\rho_{s}y\right]-F_{f}.$ (S7)
Here $F=C_{0}U_{0}(t)U_{B}/D$, since the total voltage is a sum of the bias voltage $U_{B}$ and the driving alternating voltage $U_{0}(t)$ and we only take the component that is resonant with the Helmholtz mode. The friction force $F_{f}$ can be divided into two contributions $F_{f}=F_{\mathrm{in}}+F_{\mathrm{ns}}$. The first part is the intrinsic damping of the superfluid motion, which is discussed in Section E, and can be written in the form
$F_{\mathrm{in}}=al\rho_{s}\gamma_{0}\dot{y},$ (S8)
where $\gamma_{0}$ is a dimensionless intrinsic damping factor. The second contribution is the mutual friction force between the superfluid and quantized vortices, which can be expressed Donnelly (1991) as
$\mathbf{F}_{\mathrm{ns}}=al\alpha\rho_{s}\hat{\bm{\omega}}\times[\bm{\omega}\times(\mathbf{v_{\mathrm{n}}}-\mathbf{v_{\mathrm{s}}})]+al\alpha^{\prime}\rho_{s}\bm{\omega}\times(\mathbf{v_{\mathrm{n}}}-\mathbf{v_{\mathrm{s}}}),$ (S9)
where $\mathbf{v_{\mathrm{n}}}$ and $\mathbf{v_{\mathrm{s}}}$ are velocities of the normal and the superfluid component. Macroscopic vorticity $\bm{\omega}$ represents the mean circulation of a set of parallel-oriented vortices, and $\alpha,\alpha^{\prime}$ are the temperature-dependent mutual friction coefficients Donnelly and Barenghi (1998). Due to geometrical constraints we have for the vortex density $L\approx\omega/\kappa$, where $\kappa\approx 9.97\times$10-8 m2s-1 is the quantum of circulation.
Assuming that the normal fluid component is viscously clamped to the walls, i.e. $\mathbf{v_{\mathrm{n}}}=0$, and neglecting the second non-dissipative mutual friction term, (S9) simplifies to
$F_{ns}=al\alpha\kappa\rho_{s}Lv_{\mathrm{s}}\sin^{2}{\theta},$ (S10)
where the angle $\theta$ is between $\mathbf{v_{\mathrm{s}}}$ and $\hat{\bm{\omega}}$. Assuming that the vortices are preferentially oriented perpendicularly to the superfluid flow (a consequence of 2-dimensionality), we have $\sin^{2}{\theta}=1$.
The dynamics of $L(t)$ likely occurs on a significantly longer time scale than the flow oscillation period, as evidenced by the relatively long-lived states in both the multi-stable and intermittently switching regimes (from roughly the inverse linewidth up to several seconds). In the steady state, we therefore consider $L$ to be a constant. Note, however, that this constant will depend on the RMS flow velocity amplitude (and thus drive amplitude and frequency); therefore, a Lorentzian peak shape should not be expected. A similar constancy of $L$ generated by the second sound in a 3D geometry was reported in Midlik et al. (2021).
Identifying $\dot{y}$ with $v_{\mathrm{s}}$ and substituting friction (S8),(S10) into the equation of motion (S7) results in
$\ddot{y}+(\alpha\kappa L+\gamma_{0})\dot{y}+\dfrac{2\rho_{s}ak_{p}}{l\rho^{2}(\chi V_{B}k_{p}+2A^{2})}y=\dfrac{2AC_{0}U_{B}}{\rho l(\chi V_{B}k_{p}+2A^{2})h}U_{0}(t),$ (S11)
which has the form of the equation of motion of a linear harmonic oscillator
$\ddot{y}+\tilde{\gamma}\dot{y}+\omega^{2}_{0}y=f_{0}e^{i\omega t},$ (S12)
with the total damping factor $\tilde{\gamma}$, the resonance frequency $\omega_{0}$ and a driving force $f_{0}(t)$
$\tilde{\gamma}$ $=\alpha\kappa L+\gamma_{0}=\gamma(L)+\gamma_{0},$ (S13)
$\omega_{0}^{2}$ $=\dfrac{2\rho_{s}ak_{p}}{l\rho^{2}(\chi V_{B}k_{p}+2A^{2})},$ (S14)
$f_{0}(t)$ $=\dfrac{2AC_{0}U_{B}}{\rho l(\chi V_{B}k_{p}+2A^{2})h}U_{0}(t)=f_{0}e^{i\omega t}.$ (S15)
On resonance, $\omega=\omega_{0}$ the complex response is given in the vortex-free state ($L=0$) by $y_{0}=f_{0}/(i\omega_{0}\gamma_{0})$ and for $L>0$ $y=f_{0}/(i\omega_{0}(\gamma(L)+\gamma_{0}))$, which yields
$\dfrac{y_{0}}{y}-1=\dfrac{\gamma(L)}{\gamma_{0}}.$ (S16)
Using Eq. S10 (for $\theta=\pi/2$) and the relation between the damping factor $\gamma_{0}$ and the width $\Delta_{0}$ of the un-damped resonance curve, $\gamma_{0}=2\pi\Delta_{0}$, results in
$L=\dfrac{2\pi\Delta_{0}}{\alpha\kappa}\left(\dfrac{Y_{0}}{Y}-1\right).$ (S17)
In the present work, $y_{0}$ is extrapolated from the linear regime of the force-velocity curves shown in Fig. 1D of the main text and Fig. S10.
Appendix D Calibration of force and velocity
The voltage between the device electrodes and the resulting current differ from the voltage applied to the primary winding of the center-tapped transformer and the current measured by the lock-in amplifier due to circuit losses. These losses need to be taken into account to obtain accurate velocities and driving pressure gradients. The frequency and applied voltage dependence of the circuit losses were characterized by directly measuring the oscillating voltage across the device with a separate lock-in amplifier. The current losses were characterized by injecting a known oscillating current immediately after the device on the bias side (see Fig. S1(c)). The resulting transfer functions were used to correct the driving force (proportional to the voltage across the device) and the flow velocity (proportional to the detected current).
The calculation of the mechanical quantities also requires the effective spring constant of the deformation of the substrate in the basin region. Let $C_{0}$ be the capacitance of the device corresponding to the zero DC bias voltage $U_{B}$. We characterize the effective spring constant $k_{p}$ by measuring the dependence of the device capacitance on the applied bias voltage $U_{B}$ similarly to ref. Souris et al. (2017). The force acting on the basin plates is
$F=\dfrac{CU^{2}}{2D}=k_{p}x,$ (S18)
where $D$ is the distance between the electrodes, $x$ is the mean vertical deformation of basin plates and for the capacitance $C$, we have
$C=\dfrac{A\varepsilon_{0}}{D-2x}\approx\dfrac{A\varepsilon_{0}}{D}+\dfrac{2\varepsilon_{0}A}{D^{2}}x=C_{0}+\dfrac{2C_{0}}{D}x,$ (S19)
where $A$ is the basin area, $\varepsilon$ is the vacuum permittivity, and we assumed $x\ll D$, taking into account only the first two terms of Taylor expansion. Furthermore, $x$ can be expressed from Eq. (S18) and substituting into Eq. (S19) we get
$C=C_{0}+\sigma U^{2},$ (S20)
where $\sigma=C_{0}^{2}/(D^{2}k_{p})$.
Measured capacitance includes the capacitance of the electrodes in the flow channels $C_{\mathtt{channel}}$ and a 100 nF isolation capacitor connected in series. The corrected basin capacitance is
$C_{corr}=1/(1/C_{meas}-1/100\textrm{nF})-2C_{\mathtt{channel}}.$ (S21)
The corrected $C_{corr}$ as a function of the applied voltage squared $U^{2}$ is shown in Fig. S5. The data fit with the Eq. (S20) to obtain $C_{0}$ and $k_{p}$, shown in Tab. 2
Fig.  S5: The corrected capacitance of all four Helmholtz resonators against the applied voltage. Only the DC bias voltage amplitude was changed, and the amplitude of the AC drive was kept constant at 0.02 Vrms. The black dashed line is a fit of the Eq. (S20). From the fit parameters we calculate $C_{0}$ and $k_{p}$ and these values can be found in the Tab. 2.
[…]
Corner 1.24 338.6
Rounded corner 1.19 312.9
Jet 1.45 335.7
Grid 1.62 466.4
[…]
The relevant mechanical parameters are the pressure gradient $\delta P/l$ along the flow channel resulting from the electrostatic driving force and the amplitude of the flow velocity. The equation for $\delta P/l$ can be obtained from the driving term in (S11)
$\dfrac{\delta P}{l}=\dfrac{2AC_{0}U_{B}}{l(\chi V_{B}k_{p}+2A^{2})h}U_{0},$ (S22)
where $U_{0}(t)$ was replaced by its amplitude $U_{0}$. Based on ref. Varga et al. (2020), velocity is calculated as
$v_{\mathrm{s}}=\dfrac{I}{gU_{B}C_{0}},$ (S23)
where $I$ is the detected current (corrected for losses) and $g$ is defined as
$g=\dfrac{2a\rho_{s}}{V_{B}\rho}\left(1-\dfrac{2(\varepsilon-1)}{\varepsilon}\right)\dfrac{1}{1+2\Sigma},$ (S24)
where $\varepsilon$ is the dielectric constant taken from Brooks and Donnelly (1977), and $\Sigma=\chi Dk_{p}/(2A)$.
Appendix E Fourth sound damping
Following the analysis of refs. Backhaus and Backhaus (1997); Souris et al. (2017), linear damping of the Helmholtz resonance (S8) in the temperature range 1.3 K to 2 K has two dominant contributions: viscous normal fluid motion and thermoviscous dissipation. For the inverse of resonance width $\Delta_{n}$ caused by the viscous normal flow we can write Souris et al. (2017)
$\Delta_{n}^{-1}=2\pi\dfrac{8\eta}{(D/2)^{2}}\dfrac{\rho_{s}}{\rho_{n}^{2}}\dfrac{1}{\omega_{0}^{2}},$ (S25)
where $\eta$ is the dynamic viscosity.
The second contribution is due to the heat transfer (resulting in energy loss of the oscillating flow) from the basin to the surroundings (for further details see ref. Backhaus and Backhaus (1997)). The temperature gradient between the basin and the bulk 4He is a consequence of a mechanocalorical effect, that is, the change in the superfluid density inside the basin results in a change in temperature. This effect can be characterized by the parameter
$\Gamma^{2}=\dfrac{\rho^{2}s^{2}TA^{2}(1+\Sigma)}{C_{\mathrm{th}}k_{p}},$ (S26)
where $C_{\mathrm{th}}=C_{s}V_{B}$ is the total heat capacity of 4He inside the basin ($C_{s}$ is the specific heat capacity of 4He) and $s$ is the specific entropy. For the case $\Gamma^{2}\ll 1$ (in our case $\Gamma\approx 0.22$ at 1.65 K), the influence of the additional pressure gradient caused by the presence of the temperature gradient (i.e., the fountain pressure) is negligible and for damping due to the heat flux out of the basin, it can be shown Souris et al. (2017)
$\Delta_{th}^{-1}=\dfrac{1+\Phi^{2}}{\Phi\Gamma^{2}}\left(1+\dfrac{\Gamma^{2}}{2(1+\Phi^{2})}\right)\dfrac{2\pi}{\omega_{0}},$ (S27)
where $\Phi=1/(\tau_{\mathrm{th}}\omega_{0})$. Relaxation time $\tau_{\mathrm{th}}=C_{\mathrm{th}}R_{\mathrm{th}}$ represents how quickly the system returns to thermal equilibrium. If $1/\tau_{th}<\omega_{0}$, the energy loss due to heat transport must be taken into account. The basin has two channels to exchange heat with the surroundings: helium flow through the channels or thermal conduction through the substrate. The thermal conductivity through helium in the flow channels is negligible Souris et al. (2017). The dominant limiting factor for heat conduction through the substrate is the interfacial thermal resistance (Kapitza resistance) $R_{k}$. The temperature dependence for the interface between liquid helium and aluminum is $R_{k}=28.57T^{-4.21}/A$ Souris et al. (2017), where $A$ is the basin area in cm2. Taking into account both basin plates, the total thermal resistance is $R_{\mathrm{th}}=R_{k}/2$. It should be noted, however, that the Kapitza resistance is sensitive to surface preparation, and considerable scatter exists in the reported empirical values.
The total resonance full width at half maximum (FWHM) is given by the sum of (S26) and (S27). Due to the end effects of the flow in the channels (i.e., the finite flow in the basin), the temperature oscillation inside the basin will not be uniform. Therefore, we use an effective basin area such that $A\to A/\lambda$, where $\lambda$ is an adjustable parameter. The only adjustable parameters are the scaling parameter $\lambda$ and the temperature correction (the same as for the frequency, i.e., parameter $b$ in Tab. 1). Best-fit values are listed in Tab. 1 for the four geometries.
The resulting fit to the widths of the experimentally measured resonance curves is essentially independent of the geometry of the channel, and thus Fig. S3(b) shows the average of the four fits as a single dashed line. The discrepancy between measured data and theoretical prediction is mainly in the low-temperature range; the theory underestimates the measured value by approximately a factor of two. This is likely due to uncertainty in Kapitza resistance (which is also reflected in the relatively large value of $\lambda$ in Tab. 1) since it strongly depends on the purity of the Al electrode and the structure of its surface.
Appendix F Numerical simulations
In order to better understand the behavior of the confined oscillatory flow, we performed two complementary sets of independent numerical simulations. First, a classical fluid dynamics calculation of confined incompressible turbulence based on the $k-\varepsilon$ model Davidson (2015); and second, a vortex filament simulation Hanninen and Baggaley (2014) of individual quantized vortices.
F.1 $k-\varepsilon$ simulation
The $k-\varepsilon$ is a simplified turbulence model based on the eddy viscosity approach Davidson (2015). It is a classical continuous model that does not contain any features of superfluid dynamics (e.g., mutual friction or the discreteness of superfluid vorticity), and the obtained results, therefore, have only an illustrative character. Turbulent flows of superfluid helium with large vortex densities polarized in large structures are, however, known to behave quasi-classically Babuin et al. (2014).
[…]
To obtain time-averaged behavior, 20 periods of the oscillatory flow were simulated with velocity amplitude $v_{0}=8\times 10^{-4}$ m/s and the frequency $f=\omega/(2\pi)=2000$ Hz, for the four geometries, which correspond to the typical frequency of the Helmholtz mode and velocity in the turbulent regime. Since the experiment measures behavior averaged over many oscillations, the calculated vorticity was averaged over the last period of the calculation (when the flow was already in an oscillatory steady state). The results are shown in Fig. S6. As can be seen in the cases of C, J, and G, large vortical structures are observed, which are absent in the case of R. These structures most likely correspond to the acoustic streaming patterns Duda et al. (2017). Note that the $k-\varepsilon$ model cannot simulate the inverse cascade since it assumes isotropic dissipation Davidson (2015). However, the results show that over long time periods, the acoustic flow drives a particular mode of the spatial vorticity distribution, which is consistent with the dynamical model presented in the main text.
Fig.  S6: Results of quasi 2D finite element numerical simulation in the framework of $k-\varepsilon$ turbulence model for all four chip geometries used in the experiment. We display the vorticity of fluid averaged over one period of oscillation.
F.2 Vortex filament model simulations
The motion of individual quantized vortices in uniformly oscillating superflow and stationary normal fluid confined between two walls was numerically simulated using the vortex filament model Hanninen and Baggaley (2014). Quantized vortices are represented as infinitesimally thin spatial curves, which move and deform under externally applied potential superflow and the superfluid velocity field $\mathbf{v}_{s}(\mathbf{r})$ induced by the vortices given by the Biot-Savart integral,
[…]
where $\mathbf{s}$ is a point on the vortex line, $\mathbf{s}^{\prime}$ is the unit tangent vector, $\xi$ is the arc-length measured along the vortex and the integration runs through the entire system of vortices. The equation of motion of any point $\mathbf{s}$ on the vortex is
$\frac{\partial{\mathbf{s}}}{\partial{t}}=\mathbf{v}^{\prime}_{s}+\alpha\mathbf{s}^{\prime}\times(\mathbf{v}^{\prime}_{n}-\mathbf{v}_{s})-\alpha^{\prime}\mathbf{s}^{\prime}\times\left[\mathbf{s}^{\prime}\times(\mathbf{v}^{\prime}_{n}-\mathbf{v}_{s})\right],$ (S29)
where $\alpha$ and $\alpha^{\prime}$ are temperature-dependent mutual friction coefficients Donnelly and Barenghi (1998) and $\mathbf{v}^{\prime}_{s}$ is the standard de-singularization of the Biot-Savart integral (S28) for $\mathbf{r}$ on a vortex line Hanninen and Baggaley (2014)
[…]
where $\mathbf{s}^{\prime\prime}$ is the local curvature vector, $l_{\pm}$ are the distances to forward and reverse discretization points on the vortex, and $a$ is the vortex core parameter; the integration runs over all vortices with the exception of the immediate neighborhood of $\mathbf{r}$. The vortex filament model cannot describe the nucleation of vorticity near walls; therefore, vortices are added as an initial condition. The numerical study of the experimental system with a realistic density of vortices is beyond the scope of the present work; therefore, we study the dynamics of a small number of vortices confined between two parallel walls in a doubly periodic domain in order to check the dynamical importance of three-dimensional deformations.
The vortices are discretized to a set of connected points whose spacing is dynamically maintained between $10^{-8}$ and $1.2\times 10^{-8}$ m (simulations were checked with half the discretization length and yielded identical results). The time evolution was implemented using the 4th order Runge-Kutta stepping with fixed time step $\delta t=10^{-10}$ s. Planar walls placed at $z=0$ and $z=D$ are considered to be uniformly pinning; a discretization point of a vortex attached to the wall is not allowed to move. However, the vortex can reconnect with its image on the other side of the wall, which exists because of boundary conditions of no flow normal to the wall. Depinning is handled entirely by modeling pinning and reconnections, ensuring self-consistency with the rest of the model. As a result, for a sufficiently sharp angle at the contact between the vortex and the wall, the vortex can effectively slide in a manner similar to ref Nakagawa et al. (2023).
We studied the evolution of two particular scenarios with mutual friction constants corresponding to 1.45 K, $\alpha=0.061$ and $\alpha^{\prime}=1.746\times 10^{-2}$ and $D=500$ nm – a half-loop attached to the bottom wall and a pair of reconnecting antiparallel vortex lines spanning the two walls. For all simulated cases, the imposed velocity was set to $v_{s}=0.25$ m/s along the x-axis and $v_{n}=0$. The time evolutions of the vortex configurations are shown in Fig. S7 for a pair of antiparallel vortices and Figs. S8, S9 for half-loops attached to the bottom wall oriented parallel or antiparallel with the imposed superfluid velocity $\mathbf{v}_{s}$. All initial vortex configurations studied annihilate in a few tens of microseconds with a typical flow oscillation period of 0.5 ms (see Fig. S3(a)). These results suggest that the vortex configuration will remain quasi-two-dimensional, with only transient three-dimensional structures short-lived compared to the flow oscillation period.
Fig.  S7: Typical evolution of two colliding vortices spanning the 500 nm gap between the top and bottom walls. As an initial condition, two straight vortices of opposite signs separated by 500 nm along the $y$ direction were used. The vortices are subject to a steady flow of the superfluid component with a velocity of 25 cm/s along the $x$ axis. The figure shows several snapshots leading up to the annihilation (time given by the color), with 25 ns between subsequent snapshots. The entire evolution from the initial condition to the final annihilation takes about 40 $\mu$s (the entire lead-up to the reconnection is not shown).
Fig.  S8: Annihilation of vortex half-ring with 500 nm diameter attached to the bottom wall with orientation parallel with the imposed superflow oriented along the positive x-axis direction. The initial condition was a half-circle lying in the $x=0$ plane.
Fig.  S9: As in Fig. S8 with the ring orientation anti-parallel with the imposed $\mathbf{v}_{s}$. The vortex propagates in the negative $x$ direction because its self-induced velocity exceeds the imposed superfluid velocity.
Appendix G Supplementary data
In the following, we provide an overview of the complete dataset of the measured flow velocity as a function of the driving pressure gradient (Fig. S10) or, equivalently, the vortex density as a function of the flow velocity (Figs. S11, S12, S14, and S13). For device R, Fig. S14, intermittent DC resistance in the range of several MOhm was observed across the resonator electrodes, which shunts the capacitance of the device, resulting in an unreliable calibration of force and velocity. Therefore, the comparison with the other three devices ought to be limited to qualitative features of the force-velocity, or equivalently, vortex density-velocity curves.
Figure S15 shows the decreasing lifetime of the upper state in the intermittent region of device J (complementary data to Fig. 4 in the main text), which similarly shows a power law behavior, although with a steeper critical exponent of approximately 3. Finally, Fig. S16 shows the histogram of the lower state lifetimes for three temperatures and a range of drive amplitudes corresponding to mean lifetimes of 0.1 to 0.2 s. The distribution is exponential, indicating that the switching process is memoryless. Finally, Fig. S17 shows an example of the time dependence in the intermittent switching region of device J at temperature 1.65 K.
Fig.  S10: Dependence of the superfluid velocity on the pressure gradient for all four used resonators. Both quantities are calculated from measured current and applied voltage using Eqs. (S23) and (S22). Data from different temperatures (indicated by the color) are offset for clarity by a multiplicative factor of 1.1. A, corner device; B, jet device; C, rounded corner device and D, grid device.
Fig.  S11: Vortex line density as a function of velocity for all studied temperatures (indicated in the panels) for the device C. Lines connect points from a single continuous ramp of the drive. Blue – drive is ramped up, orange – drive is ramped down.
Fig.  S12: As Fig. S11, device J.
Fig.  S13: As Fig. S11, device G.
Fig.  S14: As Fig. S11, device R. Note that the electrodes of this device were intermittently shunted by DC resistance, thus making the calibration of force and velocity unreliable.
Fig.  S15: Lifetimes of the upper state in Fig. 4 of the main text as a function of flow velocity amplitude shifted by the critical velocity $v_{c}$. The inset shows the same data in logarithmic axes against reduced velocity $(v-v_{c})/v_{c}$. The dashed line indicates a power law with exponent -3, different from the -1.33 shown by the lower state lifetimes in Fig. 4 of the main text.
Fig.  S16: Histrogram of the turbulent state lifetimes in the transitional region of device J at the three indicated temperatures for datasets with a mean lifetime in the range 0.1 to 0.2 s. A straight line in semilogarithmic axes indicates exponentially decaying dependence.
Fig.  S17: Time dependence of the observed velocity amplitude in device J in the intermittently switching region at 1.65 K (main text Fig. 4(a), middle panel), drive pressure gradient approximately 9.5 kPa/m. Maximum and minimum velocity were normalized to 0 and 1, respectively; the orange line shows the binarized signal used for the detection of transitions.